Maggies Skunk Removal Corp.s 2021 income statement listed net sales of $13.5 million, gross profit of $9.10 million, EBIT oCalculate the basic earnings power. (Round your answer to 2 decimal places.) Basic earnings power % Calculate the return on aCalculate the dividend payout. (Round your answer to 2 decimal places.) Dividend payout %
Answer #1

Given data for Maggie's Skunk Removal Corp's 2021

Net sales = $13.5 million

Gross Profit = 9.10 Million

EBIT = $6.6 million

NI = $4.2 million

Common stock dividend = $2.2 million

Total Assets = $53.5 million

equity = $22 million

Shares outstanding = 2 million

1). Gross Profit margin = Gross profit / sales = 9.1/13.5 = 67.41%

2). Operating profit margin = EBIT/Sales = 6.6/13.5 = 48.89%

3). Profit margin = NI/sales = 4.2/13.5 = 31.11%

4). Basic earning power = EBIT/Total assets = 6.6/53.5 = 12.34%

5). Return on assets = NI/Assets = 4.2/53.5 = 7.85%

6). Return on equity = NI/Equity = 4.2/22 = 19.09%

7). Dividend payout ratio = Dividend per share/earning per share or Total dividend/Net income = 2.2/4.2 = 52.38%